OpenStudy (anonymous):

Convert the polar representation of this complex number into its standard form: z=6(cos300+isin300)

OpenStudy (anonymous):

@zepdrix please help

zepdrix (zepdrix):

300 degrees is one of our special angles. It in quadrant 4.

OpenStudy (anonymous):

alright

zepdrix (zepdrix):



OpenStudy (anonymous):

mhmm

OpenStudy (anonymous):

what do i with this?

zepdrix (zepdrix):

Do you remember how to find the sine and cosine of your angles in the first quadrant? 30 45 and 60 degrees?

OpenStudy (anonymous):

no?

zepdrix (zepdrix):

Hmm that makes things kind of difficult. Haven't you learned about the unit circle and stuff? :o

OpenStudy (anonymous):

oh yeah

zepdrix (zepdrix):

You could find the reference angle for 300 degrees. Since we're in quadrant 4, we subtract our angle from 360 degrees. \(\Large\rm 360^o-\theta=360^o-300^o=60^o\) Just realize that since we're in quadrant 4, our sine (which represents the y coordinate) should be negative. \[\Large\rm \sin(300^o)=-\sin(60^o)\]Can you figure out sin(60) using a chart or something maybe? :o These are values that you should really have memorized.

OpenStudy (anonymous):

um....

OpenStudy (anonymous):

0.3048106211

zepdrix (zepdrix):

It's better if we keep it as an exact value, not a decimal approximation. http://www.sciencedigest.org/UnitCircle.gif

zepdrix (zepdrix):

If you look at the chart, at 300 degrees, the sine is giving us \(\Large\rm -\dfrac{\sqrt3}{2}\)

zepdrix (zepdrix):

How about the cosine of 300 degrees?

OpenStudy (anonymous):

5pi/3?

zepdrix (zepdrix):

No, 5pi/3 is the radian representation of 300 degrees. Cosine is the first coordinate in the set of brackets.

OpenStudy (anonymous):

1/2

zepdrix (zepdrix):

Mmmm k good.

zepdrix (zepdrix):

\[\Large\rm 6\left(\color{orangered}{\cos300^o}+\mathcal i \color{royalblue}{\sin300^o}\right)\]So plug the values in for sine,\[\Large\rm 6\left(\color{orangered}{\cos300^o}+\mathcal i \color{royalblue}{\frac{-\sqrt3}{2}}\right)\]And cosine,\[\Large\rm 6\left(\color{orangered}{\frac{1}{2}}+\mathcal i \color{royalblue}{\frac{-\sqrt3}{2}}\right)\]And then to finish it up, distribute the 6 to each term and simplify.

OpenStudy (anonymous):

so it would be 6/12 and ?

OpenStudy (anonymous):

can you please help me simplify?

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@tkhunny can you help me simplify?

zepdrix (zepdrix):

You don't know how to simplify it? 0_o

zepdrix (zepdrix):

6 times 1/2 = ....?

OpenStudy (anonymous):

6/12

OpenStudy (anonymous):

@zepdrix i dont know how to simplify the second part

zepdrix (zepdrix):

6 times 1/2 = 6/12? whut....?

zepdrix (zepdrix):

Do you remember how to multiply by a fraction? Because you're not getting the first part right. \[\Large\rm 6\cdot\frac{1}{2}=\frac{6}{1}\cdot\frac{1}{2}=\frac{6\cdot1}{1\cdot2}=\frac{6}{2}=3\]

zepdrix (zepdrix):

For the second part, same idea.\[\Large\rm 6\cdot \mathcal i \frac{(-\sqrt3)}{2}=\frac{6}{1}\cdot \mathcal i \frac{(-\sqrt3)}{2}=\mathcal i\frac{6\cdot(-\sqrt3)}{1\cdot2}=\mathcal i\frac{-6\sqrt3}{2}\]And simplify a little further. Don't do anything fancy with the i, just leave it there as part of the answer.

zepdrix (zepdrix):

I'm really surprised that you're doing this level of math and are unable to multiply fractions :( This is very worrisome.

OpenStudy (anonymous):

you really do not need to be rude, i did not mean to put a one, it was a typo, and i was very tired, but thank you for making me feel thoroughly stupid for simply pressing an extra key on my keyboard and not realizing i had done such. I worked my butt off to learn a lot of this math on my own so in all honesty, i think it's perfectly fine that i'm asking questions and what not.

zepdrix (zepdrix):

I'm very sorry Jules :( I'm still learning how to be more patient with people, espcially people who are putting in effort. And you certainly were. I'm sure you're a very special gal. I didn't mean to make you feel stupid. Sorry Jules :c Go enjoy a strawberry please.